OpenStudy (anonymous):

Is it possible for two different numbers, when squared, to give the same result? Consider the example x2=4, is x=2 the only solution? - What does this result tell you about solving an equation when the variable is squared? - How many solutions will an equation like this have? - Will there always be the same number of solutions for any equation with a squared variable? What conditions will change the number of solutions?

Nnesha (nnesha):

what is the degree of x^2 ???

Nnesha (nnesha):

degree=exponent

OpenStudy (anonymous):

I don't get what you mean ^^ I have the first 2 questions I just need this one now, - Will there always be the same number of solutions for any equation with a squared variable? What conditions will change the number of solutions? Someone please help!

Nnesha (nnesha):

highest degree tells us how many solution we going to have \[\sqrt{x^2} = \pm \sqrt{4}\] when you take square root of something your answer is going to be plus minus

OpenStudy (amistre64):

thats not quite accurate. \[\sqrt{x^2}=|x|\] and |x| is not a negative value. the issue is, that the sqrt function has been defined as having a range of [0,inf)